OpenStudy (anonymous):

how to simplify the expression -7w-21 over 7?

jimthompson5910 (jim_thompson5910):

hint: factor out 7 from the numerator

OpenStudy (anonymous):

I tried but I just don't understand the steps to take :/ Algebra 1 is sooo confusing

jimthompson5910 (jim_thompson5910):

does rewriting -7w - 21 as -w*7 - 3*7 help at all?

OpenStudy (anonymous):

it is 1w+3

OpenStudy (anonymous):

am I correct?

jimthompson5910 (jim_thompson5910):

close, but you factor out 7 to get 7( -w - 3)

OpenStudy (anonymous):

so it would be -7w-21?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

i factored out the 7 for a specific reason

OpenStudy (anonymous):

I'm thinking it would make it easier.....

jimthompson5910 (jim_thompson5910):

it's so the 7's will cancel

OpenStudy (anonymous):

\[\frac{-7w-21}{7}=\frac{-7w}{7}+\frac{-21}{7}=-w+-3=-w-3 \]